OpenStudy (anonymous):

if p is the subject of this equation 4p-3y/2=m/2

OpenStudy (anonymous):

I do not understand the question.

OpenStudy (anonymous):

make p the subject of this equation: 4p-3y/2 = m/2

OpenStudy (anonymous):

cancelling 2 from both sides ===> 4p-3y = m ===> 4p= m+3y ==> p = (1/4)(m+3y)